# Generalized Frobenius Manifold Structures on the Orbit Spaces of Affine Weyl Groups I

### v1

@@ 4.3. The Monodromy group of $\mathcal{M}(R,\omega)$
Note that $z^{1},\cdots,z^{\ell}$ are coordinates on $\mathcal{M}\setminus(D\cup\Sigma^{\prime})$, so intuitively speaking, $r(\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime})))$ consists of those isometric transformations on $(x^{1}\bigr{|}_{\lambda=0},\cdots,x^{\ell}\bigr{|}_{\lambda=0})$ that preserve $(z^{1},\cdots,z^{\ell})$. Similarly, $r(\pi_{1}(\overline{\mathcal{M}}\setminus\overline{\Sigma}))$ consists of those isometric transformations on $(x^{1}\bigr{|}_{\lambda=0},\cdots,x^{\ell}\bigr{|}_{\lambda=0})$ that preserve $(t^{1},\cdots,t^{\ell})$. From Theorem 4.2 we know that, if $t^{1},\cdots,t^{\ell}$ are preserved under some transformations, so are $z^{1},\cdots,z^{\ell}$. This fact implies the relation (4.39).
We are ready to present the following theorem.
Theorem 4.6.
For the generalized Frobenius manifold ¯$\overline{\mathcal{M}}=\mathcal{M}(R,\omega)$, the group $r(\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime})))$ is isomorphic to a subgroup of $W_{a}(R)\cong W(R)\rtimes\mathbb{Z}^{\ell}$. More precisely,
$r(\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime})))\cong\mathrm{Stab}(\omega)\rtimes\mathbb{Z}^{\ell}.$ (4.48)
Here $\mathrm{Stab}(\omega)$ is the stabilizer of $\omega$ in $W(R)$.
Proof.
Due to Proposition 4.5, we only need to determine all the isometric actions on $x^{i}\bigr{|}_{\lambda=0}$ that preserve $\mathcal{M}\setminus(D\cup\Sigma^{\prime})$, i.e., preserve $z^{1},\cdots,z^{\ell}$. From (2.31) and (2.41) it follows that
$z^{j}\bigr{|}_{\lambda=0}=\frac{1}{N_{j}}\sum_{\sigma\in H_{j}}\mathrm{e}^{2\pi i(\sigma^{*}(\omega_{j}),\hat{\mathbf{x}})}.$ (4.49)
Here
$H_{j}=\{\sigma\in W(R)\mid(\omega_{j}-\sigma^{*}(\omega_{j}),\omega)=0\}$
is a subgroup of $W(R)$. For any given element $\tilde{\phi}\in\mathrm{Iso}(E^{\ell})$, we decompose it into linear and the affine parts as follows:
$\tilde{\phi}(\mathbf{t})=\phi(\mathbf{t})+v^{1}\alpha_{1}^{\vee}+\dots+v^{\ell}\alpha_{\ell}^{\vee}.$ (4.50)
It acts on the the Fourier polynomials $z^{j}$ by
$\tilde{\phi}^{*}z^{j}\big{|}_{\lambda=0}$ $=\frac{1}{N_{j}}\sum_{\sigma\in H_{j}}\mathrm{e}^{2\pi i(\sigma^{*}(\omega_{j}),\tilde{\phi}(\hat{\mathbf{x}}))}$ (4.51)
$=\frac{1}{N_{j}}\sum_{\sigma\in H_{j}}\mathrm{e}^{2\pi i(\sigma^{*}(\omega_{j}),\phi(\hat{\mathbf{x}})+v^{1}\alpha_{1}^{\vee}+\dots+v^{\ell}\alpha_{\ell}^{\vee})},$ (4.52)
so it preserves $z^{j}\bigr{|}_{\lambda=0}$, $j=1,\cdots,\ell$ if and only if $\phi$ induces a permutation on $H_{j}$ by right action $\sigma\to\sigma\phi$ and
$(\sigma^{*}(\omega_{j}),v^{1}\alpha_{1}^{\vee}+\cdots+v^{\ell}\alpha_{\ell}^{\vee})$
is always an integer for each $\sigma\in H_{j}$, $j=1,\cdots,\ell$. These conditions are equivalent to
$\phi\in\bigcap_{j=1}^{\ell}H_{j},\quad v^{1},\cdots,v^{\ell}\in\mathbb{Z}.$ (4.53)
This implies that
$r(\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime})))=\bigcap_{j=1}^{\ell}H_{j}\rtimes\mathbb{Z}^{\ell}.$ (4.54)
Now let us compute $\bigcap_{j=1}^{\ell}H_{j}$. Note that $\sigma\in H_{j}$, $j=1,\cdots,\ell$ if and only if
$(\omega_{j},\omega)=(\omega_{j},\sigma(\omega)),\quad j=1,\cdots,\ell.$ (4.55)
Since $\omega_{j},j=1,\cdots,\ell$ span $V\otimes\mathbb{C}$, (4.55) is equivalent to $\sigma(\omega)=\omega$. Therefore,
$\bigcap_{j=1}^{\ell}H_{j}=\mathrm{Stab}(\omega).$ (4.56)
It follows from (4.54) and (4.56) that $r(\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime})))=\mathrm{Stab}(\omega)\rtimes\mathbb{Z}^{\ell}$. The theorem is proved. ∎
By using (4.39), we have the following corollary.
Corollary 4.7.
The monodromy group of the generalized Frobenius manifold $\mathcal{M}(R,\omega)$ is isomorphic to a subgroup of $\mathrm{Stab}(\omega)\rtimes\mathbb{Z}^{\ell}$.
It comes out a natural question to ask: is $\mathrm{Mono}(\mathcal{M}(R,\omega))$ exactly isomorphic to $\mathrm{Stab}(\omega)\rtimes\mathbb{Z}^{\ell}$? Unfortunately, the answer to this question is negative, and we will see counterexamples in Sec. 5. However, in some special cases, the isomorphism holds true. One simple case is when ¯$\overline{\mathcal{M}}=\mathcal{M}\setminus D$. In this case we have ¯$\overline{\mathcal{M}}\setminus\overline{\Sigma}\cong(\mathcal{M}\setminus D)\setminus\Sigma^{\prime}$, hence
$\mathrm{Mono}(\overline{\mathcal{M}})=r(\pi_{1}(\overline{\mathcal{M}}\setminus\overline{\Sigma}))=r(\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime})))~{}\cong~{}\mathrm{Stab}(\omega)\rtimes\mathbb{Z}^{\ell}.$ (4.57)
We will also see examples which belong to this case in Sec. 5.

### v3

@@ 4.3. The Monodromy group of $\mathcal{M}(R,\omega)$
Now we are ready to compute the monodromy group $\mathrm{Mono}(\mathcal{M}_{D})$.
Theorem 4.11.
For the generalized Frobenius manifold $\mathcal{M}_{D}=\mathcal{M}(R,\omega)$, its monodromy group $\mathrm{Mono}(\mathcal{M}_{D})$ is isomorphic to a subgroup of $W_{a}(R)\cong W(R)\ltimes\mathbb{Z}^{\ell}$. More precisely,
$\mathrm{Mono}(\mathcal{M}_{D})\cong\mathrm{Stab}_{W}(\omega)\ltimes\mathbb{Z}^{\ell}.$ (4.50)
Here $\mathrm{Stab}_{W}(\omega)$ stands for the stabilizer of $\omega$ in $W(R)$.
We first prove three lemmas.
Lemma 4.12.
Let $\mathbb{F}\hookrightarrow\mathbb{E}$ be a field extension, $x_{1},\dots,x_{m}\in\mathbb{E}$ be some nonzero algebraic elements over $\mathbb{F}$. Suppose $a_{1},\dots,a_{m}\in\mathbb{E}$ such that $a_{1}x_{1}^{n}+\cdots+a_{m}x_{m}^{n}\in\mathbb{F}$ holds for any $n\in\mathbb{Z}_{>0}$, then it also holds for $n=0$, i.e., we have $a_{1}+\cdots+a_{m}\in\mathbb{F}$.
Proof.
Since $x_{i}\not=0$, the minimal polynomial $p_{i}$ of $x_{i}$ over $\mathbb{F}$ has a nonzero constant term (even if $x_{i}\in\mathbb{F}$). Let $p=p_{1}\cdots p_{m}$, then $a_{1}p(x_{1})+\cdots+a_{m}p(x_{m})=0$. Since $p$ has a nonzero constant term, we have $a_{1}+\cdots+a_{m}\in\mathbb{F}$. ∎
Lemma 4.13.
For any $\rho\in\bigoplus_{j=1}^{\ell}\mathbb{Z}\omega_{j}$, we have
$Q_{\rho}(\hat{\mathbf{x}})=\sum_{\sigma\in\mathrm{Stab}_{W}(\omega)}\mathrm{e}^{2\pi i(\rho,\sigma(\hat{\mathbf{x}}))}\in\mathbb{C}\left(z^{1}\big|_{\lambda=0},\dots,z^{\ell}\big|_{\lambda=0}\right).$ (4.51)
Proof.
For any $\rho\in\bigoplus_{j=1}^{\ell}\mathbb{Z}\omega_{j}$, we denote
$\rho=\rho_{1}\omega_{1}+\cdots+\rho_{\ell}\omega_{\ell},\quad\rho_{j}\in\mathbb{Z}.$ (4.52)
We first consider the case when each $\rho_{j}>0$. In what follows, we briefly call this case by $\rho>0$. Note that
$\sum_{\sigma\in W(R)}\mathrm{e}^{2\pi i(\rho,\sigma(\mathbf{x})-c\omega)}=\sum_{\sigma\in W(R)}\mathrm{e}^{2\pi i(\rho,\sigma(\hat{\mathbf{x}})+c\sigma(\omega)-c\omega)}\in\mathscr{A}^{W}[\lambda^{\varepsilon/\kappa}],$ (4.53)
where $\varepsilon=\mathrm{gcd}\{(\omega_{r},\alpha_{r})\mid r=1,\dots,\ell\}$. Indeed, (4.53) is $W_{a}(R)$-invariant, and equation (2.18) implies that
$\sigma(\omega)-\omega=-\sum_{r=1}^{\ell}n_{r}\alpha_{r}$
for some $n_{r}\in\mathbb{Z}_{\geq 0}$, which indicates that $(\rho,\sigma(\omega)-\omega)\in\varepsilon\mathbb{Z}_{\leq 0}$. Since $\mathscr{A}^{W}=\mathbb{C}[z^{1},\dots,z^{\ell};\lambda]$, by putting $\lambda=0$ in (4.53), we obtain
$\sum_{\begin{subarray}{c}\sigma\in W(R)\\ (\rho,\sigma(\omega)-\omega)=0\end{subarray}}\mathrm{e}^{2\pi i(\rho,\sigma(\hat{\mathbf{x}}))}\in\mathbb{C}\left[z^{1}\big|_{\lambda=0},\dots,z^{\ell}\big|_{\lambda=0}\right].$
Therefore, we only need to show that for any $\sigma\in W(R)$, $(\rho,\sigma(\omega)-\omega)=0$ implies $\sigma(\omega)=\omega$. In fact, if $(\rho,\sigma(\omega)-\omega)=0$, we have
$(\rho,\sigma(\omega)-\omega)=(\rho,-\sum_{r=1}^{\ell}n_{r}\alpha_{r})=\sum_{r=1}^{\ell}\rho_{r}n_{r}(\omega_{r},\alpha_{r})=0.$
Since $\rho_{r}>0$, $n_{r}\geq 0$ and $(\omega_{r},\alpha_{r})>0$, $r=1,\dots,\ell$, we have $n_{r}=0$, $r=1,\dots,\ell$, hence $\sigma(\omega)=\omega$. Therefore, $Q_{\rho}\in\mathbb{C}(z^{1}\big|_{\lambda=0},\dots,z^{\ell}\big|_{\lambda=0})$ for any $\rho>0$.
In particular, we know that for any $\rho>0$ and $\sigma\in\mathrm{Stab}_{W}(\omega)$, $\mathrm{e}^{2\pi i(\rho,\sigma(\hat{\mathbf{x}}))}$ is algebraic over $\mathbb{C}(z^{1}\big|_{\lambda=0},\dots,z^{\ell}\big|_{\lambda=0})$. In fact, since for any $n\in\mathbb{Z}_{>0}$, we have $n\rho>0$, hence
$Q_{n\rho}(\hat{\mathbf{x}})=\sum_{\sigma\in\mathrm{Stab}_{W}(\omega)}\mathrm{e}^{2\pi i(n\rho,\sigma(\hat{\mathbf{x}}))}\in\mathbb{C}\left(z^{1}\big|_{\lambda=0},\dots,z^{\ell}\big|_{\lambda=0}\right).$
In other words, all the Newton sums of $\{\mathrm{e}^{2\pi i(\rho,\sigma(\hat{\mathbf{x}}))}\mid\sigma\in\mathrm{Stab}_{W}(\omega)\}$ belong to the field $\mathbb{C}\left(z^{1}\big|_{\lambda=0},\dots,z^{\ell}\big|_{\lambda=0}\right)$, hence so are their symmetric polynomials
$r_{n}=\sum_{\begin{subarray}{c}I\subseteq\mathrm{Stab}_{W}(\omega)\\ \#I=n\end{subarray}}~\prod_{\sigma\in I}\mathrm{e}^{2\pi i(\rho,\sigma(\hat{\mathbf{x}}))}\in\mathbb{C}\left(z^{1}\big|_{\lambda=0},\dots,z^{\ell}\big|_{\lambda=0}\right).$
This implies that they are algebraic over $\mathbb{C}\left(z^{1}\big|_{\lambda=0},\dots,z^{\ell}\big|_{\lambda=0}\right)$.
Now let us consider the general case. For any $\rho\in\bigoplus_{j=1}^{\ell}\mathbb{Z}\omega_{j}$, we take $\xi\in\bigoplus_{j=1}^{\ell}\mathbb{Z}\omega_{j}$ and $\xi>0$ such that $\rho+\xi>0$. Then for any $n\in\mathbb{Z}_{>0}$, we have $\rho+n\xi>0$, hence
$Q_{\rho+n\xi}(\hat{\mathbf{x}})=\sum_{\sigma\in\mathrm{Stab}_{W}(\omega)}\mathrm{e}^{2\pi i(\rho+n\xi,\sigma(\hat{\mathbf{x}}))}\in\mathbb{C}\left(z^{1}\big|_{\lambda=0},\dots,z^{\ell}\big|_{\lambda=0}\right).$
Since $\mathrm{e}^{2\pi i(\xi,\sigma(\hat{\mathbf{x}}))}$ is algebraic over $\mathbb{C}\left(z^{1}\big|_{\lambda=0},\dots,z^{\ell}\big|_{\lambda=0}\right)$, it follows from Lemma 4.12 that
$Q_{\rho}(\hat{\mathbf{x}})=\sum_{\sigma\in\mathrm{Stab}_{W}(\omega)}\mathrm{e}^{2\pi i(\rho,\sigma(\hat{\mathbf{x}}))}\in\mathbb{C}\left(z^{1}\big|_{\lambda=0},\dots,z^{\ell}\big|_{\lambda=0}\right).$
Therefore, we have proved the lemma. ∎
Lemma 4.14.
Let $R$ be a root system, $\omega$ be a weight and $\mathscr{A}$ be the corresponding $\lambda$-Fourier polynomial ring. Denote by $\mathscr{K}=\mathrm{Frac}(\mathscr{A})$ the fraction field of $\mathscr{A}$, then $W(R)$ acts on $\mathscr{K}$, and
$\mathrm{Inv}_{\mathscr{K}}\left(\mathrm{Stab}_{W}(\omega)\right)=\mathbb{C}\left(z^{1}\big|_{\lambda=0},\dots,z^{\ell}\big|_{\lambda=0};\lambda\right).$
Here $\mathrm{Inv}_{\mathscr{K}}(\mathrm{Stab}_{W}(\omega))$ refers to the fixed subfield of $\mathscr{K}$ with respect to the action of $\mathrm{Stab}_{W}(\omega)$. In particular, the field extension $\mathbb{C}(z^{1}\big|_{\lambda=0},\dots,z^{\ell}\big|_{\lambda=0};\lambda)\hookrightarrow\mathscr{K}$ is Galois.
Proof.
From (2.31) and (2.36) we know that
$z^{j}\bigr|_{\lambda=0}=\frac{1}{N_{j}}\sum_{\sigma\in H_{j}}\mathrm{e}^{2\pi i(\sigma^{-1}(\omega_{j}),\hat{\mathbf{x}})}$, (4.54)
where
$H_{j}=\{\sigma\in W(R)\mid(\omega_{j},\omega-\sigma(\omega))=0\}.$ (4.55)
Since for any $\phi\in\mathrm{Stab}_{W}(\omega)$, $\sigma\mapsto\sigma\phi$ induces a permutation on $H_{j}$, we have
$z^{j}\big|_{\lambda=0}\in\mathrm{Inv}_{\mathscr{K}}(\mathrm{Stab}_{W}(\omega)).$
Thus we only need to prove that
$\mathrm{Inv}_{\mathscr{K}}(\mathrm{Stab}_{W}(\omega))\subseteq\mathbb{C}\left(z^{1}\big|_{\lambda=0},\dots,z^{\ell}\big|_{\lambda=0};\lambda\right).$
We first show that
$\mathbb{C}[\mathrm{e}^{\pm 2\pi ix^{j}}\mid j=1,\dots,\ell]^{\mathrm{Stab}_{W}(\omega)}\subseteq\mathbb{C}\left(z^{1}\big|_{\lambda=0},\dots,z^{\ell}\big|_{\lambda=0}\right).$ (4.56)
In fact, for any $q\in\mathbb{C}[\mathrm{e}^{\pm 2\pi ix^{j}}\mid j=1,\dots,\ell]^{\mathrm{Stab}_{W}(\omega)}$, we denote
$q=\sum_{\rho\in\bigoplus_{j=1}^{\ell}\mathbb{Z}\omega_{j}}A_{\rho}\mathrm{e}^{2\pi i(\rho,\hat{\mathbf{x}})}$,
where only finitely many $A_{\rho}$ is not zero. Since $q$ is $\mathrm{Stab}_{W}(\omega)$-invariant, we have
[…]
By using Lemma 4.13, we arrive at (4.56).
We next show that
$\mathscr{A}^{\mathrm{Stab}_{W}(\omega)}\subseteq\mathbb{C}(z^{1}\big|_{\lambda=0},\dots,z^{\ell}\big|_{\lambda=0};\lambda).$ (4.57)
Recall that $W(R)$ does not act on $\mathscr{A}$, since its action may introduce terms with negative power of $\lambda$. However, its subgroup $\mathrm{Stab}_{W}(\omega)$ acts on $\mathscr{A}$. In fact, for each $q\in\mathscr{A}$, we can write it as
$q(\hat{\mathbf{x}};\lambda)=\sum_{r}\lambda^{r}q_{r}(\hat{\mathbf{x}}),\quad q_{r}\in\mathbb{C}[\mathrm{e}^{\pm 2\pi ix^{j}}\mid j=1,\dots,\ell].$
Then an element $\phi\in\mathrm{Stab}_{W}(\omega)$ acts on $q$ by
$),\displaystyle(\phi^{*}q)(\hat{\mathbf{x}};\lambda)=\sum_{r}\lambda^{r}q_{r}(\phi(\mathbf{x})-c\omega)=\sum_{r}\lambda^{r}q_{r}(\phi(\hat{\mathbf{x}}$)), (4.58)
where we used the fact that $\phi(\omega)=\omega$. Therefore, $\mathrm{Stab}_{W}(\omega)$ indeed acts on $\mathscr{A}$. Note that (4.58) also implies that if $q\in\mathscr{A}^{\mathrm{Stab}_{W}(\omega)}$, then each $q_{r}$ is also $\mathrm{Stab}_{W}(\omega)$-invariant, hence it follows from (4.56) that
$q_{r}\in\mathbb{C}(z^{1}\big|_{\lambda=0},\dots,z^{\ell}\big|_{\lambda=0}),$
and $q\in\mathbb{C}(z^{1}\big|_{\lambda=0},\dots,z^{\ell}\big|_{\lambda=0};\lambda)$.
Finally we show that for any $p/q\in\mathrm{Inv}_{\mathscr{K}}(\mathrm{Stab}_{W}(\omega))$, where $p,q\in\mathscr{A}$, we have
$\frac{p}{q}\in\mathbb{C}(z^{1}\big|_{\lambda=0},\dots,z^{\ell}\big|_{\lambda=0};\lambda).$
In fact, we have
$\frac{p}{q}=\frac{p\cdot\prod_{\phi\in\mathrm{Stab}_{W}(\omega)\setminus\{\mathrm{id}\}}\phi^{*}q}{\prod_{\phi\in\mathrm{Stab}_{W}(\omega)}\phi^{*}q}=:\frac{\tilde{p}}{\tilde{q}}$.
Here $\tilde{q}\in\mathscr{A}^{\mathrm{Stab}_{W}(\omega)}$, hence $\tilde{p}\in\mathscr{A}^{\mathrm{Stab}_{W}(\omega)}$. Therefore, by using (4.57) we have
$\frac{p}{q}\in\mathbb{C}(z^{1}\big|_{\lambda=0},\dots,z^{\ell}\big|_{\lambda=0};\lambda).$
The lemma is proved. ∎
Proof of Theorem 4.11.
Due to Proposition 4.10, we only need to determine all the isometric actions on $V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma)$ that preserve $\mathcal{M}_{D}\setminus\Sigma$, i.e. actions on coordinates $x^{i}$ that preserves $z^{j}\bigr|_{\lambda=0}$. Recall that
$z^{j}\bigr|_{\lambda=0}=\frac{1}{N_{j}}\sum_{\sigma\in H_{j}}\mathrm{e}^{2\pi i(\sigma^{-1}(\omega_{j}),\hat{\mathbf{x}})}$. (4.59)
For any given element $\tilde{\phi}\in\mathrm{Iso}(V\otimes\mathbb{C})$, we decompose it into linear and the affine parts as follows:
$\tilde{\phi}(\hat{\mathbf{x}})=\phi(\hat{\mathbf{x}})+v^{1}\alpha_{1}^{\vee}+\dots+v^{\ell}\alpha_{\ell}^{\vee}.$ (4.60)
Then $\tilde{\phi}$ acts on the the Fourier polynomials $z^{j}$ by
$(\tilde{\phi}^{*}z^{j})\big|_{\lambda=0}$ $=\frac{1}{N_{j}}\sum_{\sigma\in H_{j}}\mathrm{e}^{2\pi i(\sigma^{-1}(\omega_{j}),\tilde{\phi}(\hat{\mathbf{x}}))}$ (4.61)
$=\frac{1}{N_{j}}\sum_{\sigma\in H_{j}}\mathrm{e}^{2\pi i(\sigma^{-1}(\omega_{j}),\phi(\hat{\mathbf{x}})+v^{1}\alpha_{1}^{\vee}+\dots+v^{\ell}\alpha_{\ell}^{\vee})},$ (4.62)
where $H_{j}$ is defined in (4.55). Therefore, the action preserves all $z^{j}\bigr|_{\lambda=0}$, $j=1,\dots,\ell$ if and only if both of the following conditions hold:
  1. (1)
    For each $j=1,\dots,\ell$, $\phi$ induces a permutation on $\{\sigma^{-1}(\omega_{j})\mid\sigma\in H_{j}\}$ by right action $\sigma\mapsto\sigma\phi$.
  2. (2)
    for each $j=1,\dots,\ell$, $\sigma\in H_{j}$,
    $(\sigma^{-1}(\omega_{j}),v^{1}\alpha_{1}^{\vee}+\cdots+v^{\ell}\alpha_{\ell}^{\vee})$
    is an integer.
We now proceed to respectively show that condition (1) is equivalent to $\phi\in\mathrm{Stab}_{W}(\omega)$, and condition (2) is equivalent to $v^{1},\dots,v^{\ell}\in\mathbb{Z}$.
Suppose $\phi\in\mathrm{Stab}_{W}(\omega)$, then $\sigma\mapsto\sigma\phi$ induces a permutation on $H_{j}$, hence on $\{\sigma^{-1}(\omega_{j})\mid\sigma\in H_{j}\}$. On the other hand, if condition (1) holds, we are to show that $\phi\in\mathrm{Stab}_{W}(\omega)$. We consider the canonical action of $\phi$ on $\mathscr{K}$ given by $\phi^{*}q(\mathbf{x},\lambda)=q(\phi(\mathbf{x}),\lambda)$. According to Lemma 4.14, we know that
$\mathrm{Gal}\left(\mathscr{K}/\mathbb{C}\left(z^{1}|_{\lambda=0},\dots,z^{\ell}|_{\lambda=0};\lambda\right)\right)=\mathrm{Stab}_{W}(\omega).$
Condition (1) implies that $\phi^{*}$ fixes $\mathbb{C}\left(z^{1}|_{\lambda=0},\dots,z^{\ell}|_{\lambda=0};\lambda\right)$, hence we only need to show that $\phi^{*}(\mathscr{K})\subseteq\mathscr{K}$. Note that $\mathscr{K}$ is generated by $\mathrm{e}^{2\pi ix_{1}},\dots,\mathrm{e}^{2\pi ix_{\ell}}$ and $\lambda$ as $\mathbb{C}$-algebra, and we have $\phi^{*}\lambda=\lambda$, it suffices to show that $\phi^{*}\mathrm{e}^{2\pi ix_{1}}\in\mathscr{K}$. In fact, by considering $\mathrm{id}\in H_{j}$, $j=1,\dots,\ell$, we know from condition (1) that
$\phi^{*}(\omega_{j})=\sigma_{j}(\omega_{j}),\quad\exists~\sigma_{j}\in H_{j}.$
Then in particular, for each $j=1,\dots,\ell$, we have
$\phi^{*}\mathrm{e}^{2\pi ix^{j}}=\phi^{*}\mathrm{e}^{2\pi i(\omega_{j},\mathbf{x}-c\omega)}=\mathrm{e}^{2\pi i(\omega_{j},\phi(\mathbf{x})-c\omega)}=\mathrm{e}^{2\pi i(\sigma_{j}^{-1}(\omega_{j}),\hat{\mathbf{x}})}$.
Here we used $(\sigma_{j}^{-1}(\omega_{j})-\omega_{j},\omega)=0$. From (2.20) we know that for each $r=1,\dots,\ell$, $(\sigma_{j}^{-1}(\omega_{j}),\alpha_{r}^{\vee})$ is an integer, so
$\phi^{*}\mathrm{e}^{2\pi ix^{j}}=\mathrm{e}^{2\pi i\sum_{r}x^{r}(\sigma_{j}^{-1}(\omega_{j}),\alpha_{r}^{\vee})}\in\mathscr{A},\quad j=1,\dots,\ell.$
It follows immediately that $\phi^{*}(\mathscr{K})\subseteq\mathscr{K}$. Hence $\phi\in\mathrm{Stab}_{W}(\omega)$.
Suppose $v^{1},\dots,v^{\ell}\in\mathbb{Z}$, then for any $\sigma\in H_{j}$, we know from (2.20) that for each $i=1,\dots,\ell$ $(\sigma^{-1}(\omega_{j}),\alpha_{i}^{\vee})$ is an integer, so is their $\mathbb{Z}$-combination $(\sigma^{-1}(\omega_{j}),v^{1}\alpha_{1}^{\vee}+\cdots+v^{\ell}\alpha_{\ell}^{\vee})$. On the other hand, if condition (2) holds, then for any $j=1,\dots,\ell$, by considering $\mathrm{id}\in H_{j}$, we obtain that $v^{j}\in\mathbb{Z}$.
Therefore, we have $\mathrm{Mono}(\mathcal{M}_{D})=\mathrm{Stab}_{W}(\omega)\ltimes\mathbb{Z}^{\ell}$. ∎
By using Proposition 4.8, we have the following corollary.
Corollary 4.15.
The monodromy group of $\widehat{\mathcal{M}}_{D}$ is isomorphic to a subgroup of $\mathrm{Stab}_{W}(\omega)\ltimes\mathbb{Z}^{\ell}$.
In general, the monodromy group of $\widehat{\mathcal{M}}_{D}$ is not isomorphic to that of $\mathcal{M}_{D}$, and we will give a counterexample in Sect. 5. However, in some cases we have $\widehat{\mathcal{M}}_{D}=\mathcal{M}_{D}$, so they have isomorphic monodromy groups. We will also see such examples in Sect. 5.
@@
The group $\mathrm{Stab}_{W}(\omega)$ is a parabolic subgroup of $W(R)$ [6]. Explicitly, we have
$\mathrm{Stab}_{W}(\omega)=W_{S^{c}}=\langle\sigma_{i}\mid i\in S^{c}\rangle,$ (4.63)
where $S$ is given by (2.29) and $S^{c}$ is its complement in $\{1,\cdots,\ell\}$.